ROTATIONAL SPECTRA OF THE H2-H2O AND H2-D2O CLUSTERS OBSERVED BY MILLIMETER-WAVE SPECTROSCOPY COMBINED WITH A PULSED SUPERSONIC JET TECHNIQUE

C. J. WHITHAM, \underline{KEIICHI TANAKA}, and K. HARADA, Department of Chemistry, Kyushu University, Fukuoka, 812-8581, Japan.

Millimeter-wave absorption spectroscopy combined with a pulsed-jet expansion technique was applied to the measurement of the rotational transitions of of the(o)H_{2}-H_{2}O and (o)H_{2}-D_{2}O complexes in the frequency region up to 400 GHz. Rotational transitions of the complexes for two internal rotation states of the water part, Σ(0_{00}) and Π(1_{01}), were observed. The H_{2} part was, however, limited to the ortho species with nuclear spin of one and internal angular moment of j_{H2} = 1 and k_{H2} = 0. No spectral lines for the para-H_{2} species have been detected. The Π(1_{01}) state is split into doublet (e and f), due to the Coriolis interaction between the Π and Σ sublevels in the (1_{01}) state. Analysis of the Coriolis interaction give us the anistropy of potential energy surface around H_{2}O, and the Σ sublevel is estimated to be located about 10 cm^{−1} above the Π sublevel. The center of mass bond length between H_{2} and water was derived to be 3.598 and 3.610 A, for (o)H_{2}-H_{2}O and (o)H_{2}-D_{2}O, respectively, from the observed rotational constants of the Σ(0_{00}) state. Simple potential fuction was determined by fitting to the observed rotational lines.